Polonium is the only metal that forms a simple cubic crystal structure. Use the fact that the density of polonium is to calculate its atomic radius (in ).


pm


1
Expert's answer
2021-11-05T00:56:48-0400

edge length = a

molar mass of polonium = 209 g/mol

number of atoms = Z = 1

density = Z M / N a^3

9.32 = 1 x 209 / (6.023 x 10^23 x a^3)

a^3 = 3.72 x 10^-23

a = 3.34 x 10^-8 cm

in simple cubic radius and edge length relation :

a = 2r

r = a/2 = 3.34 x 10^-8 / 2

r = 1.67 x 10^-8 cm

radius of polonium = 167 pm

density = 8.9 g/cm^3